The depletion of ozone (O3) in the stratosphere has been a matter of great concern among scientists in recent years. It is believed that ozone can react with nitric oxide (NO) that is discharged from the high altitude jet plane, the SST. The reaction is


O3+NO product> O2+NO2


If 0.740 g of O3 reacts with 0.670 g of NO, how many grams of NO2 will be produced? Which compound is the limiting reagent? Calculate the number of moles of the excess reagent remaining at the end of the reaction.


1
Expert's answer
2021-10-08T02:07:08-0400

Molar mass of O3= 48

0.74/48= 0.0154

Molar mass of NO= 30

0.67/30= 0.022

Molar mass of NO2= 46

= 0.0154+0.022= 0.0376moles